Amber Company produces iron table and chair sets. During October, Ambers costs were as follows: Actual purchase price Actual4. Compute the standard direct labor rate for October. (Round your answer to 2 decimal places.) Standard Direct Labor Rate 5.
Answer #1

1) Total cost of purchase = total quantity of sets purchased × Actual Purchase price per set

= 1015000 lbs ×$1.3 per lb

= $ 1319,500

2) Direct Material Price variance= Actual quantity( Std price- Actual price)

= 1015000(1.10-1.30)

= $ 203000(Adverse)

3) Direct material quantity variance = Std price(std quantity- actual quantity)

= 1.10(870000-1015000)

=$ 159500(Adverse)

4) Standard labour hour rate=

Direct labour rate variance = $4500

(std rate - actual rate)Actual labour hours = $4500

std rate - 6.50 = 4500/7000= 0.6428

Std rate= 6.50+ 0.6428

=$ 7.1428 per hour

5) Direct labour efficiency variance= (std hours-Actual hours)× std labour rate

= $15000- $7000(7.1428)

= $ 57142.40(favourable)